# Premerger phenomena in neutron-star binary coalescences

### v1

@@ V.2.1 Excluded events
The first two of these are described by Troja et al. 398, and are remarkable in that the waiting times exceed $\sim 100$ s. The authors remark several times however that they “are unable to confidently determine whether they are real features” as they were not seen by instruments other than Swift and are of relatively low significance. Although the possibility of such a long waiting time could offer exciting insight into neutron-star structure (cf. Fig. 9), we do not comment on these events further here. The final BATSE event is described by Koshut et al. 215, also with a very long waiting time of $\sim 75$ s. The significance is, however, similarly unclear (as is perhaps visually evident from the fourth-to-last panel seen in their figure 3). Note that Wang et al. 420 found a precursor from 080702A but much closer to the main event; it could be, therefore, that this is the second example of a known double precursor event after 090510 (see Sec. V.6.5). In the case of GRB 0911117, Troja et al. 398 find a significance of only $\sim 1.8\sigma$ and Zhong et al. 443 remark that the data can no longer be downloaded from the Swift archives (although this event may have also been recorded by Suzaku). We exclude this otherwise unremarkable event also (waiting time $\sim 2.7$ s).
@@ V.3 Rarity
Uncertainties notwithstanding with respect to the long-short classification issues detailed above, classifications for merger-driven precursors also depend on a variety of systematics. For instance, some authors require that a genuine precursor must precede the main burst by more than the burst $T_{90}$ [e.g. 398, 270]. This would mean there is a clear separation of timescales (as in Fig. 13), and not a possible blur period where it is unclear whether the precursor was related to burst ramp-up (cf. GRB 150922A). Some other studies allow for arbitrarily short waiting times relative to the main burst for early flashes to qualify for precursor status [see 443, 96, 420, for discussions].
@@
Another more obvious issue relates to sample size and instruments, with some events only being seen by Fermi, and/or Swift, and/or INTEGRAL [some events even by Agile, KONUS, Suzaku, Messenger, and HEND-Odyssey; 270, cf. Table 3]. As such one has to make a statistical decision about the significance required to define a precursor with respect to one or more instruments. Taking a high cut-off for some information criterion (e.g. Akaike or Bayesian) will naturally reduce the event rate and hence predict that precursors are rarer. As noted by Suvorov and Kokkotas 378 and others, even these deceptively small systematics can change event rate predictions from $\sim 0.4\%$ [270] to $\sim 10\%$ [398], with the most recent studies falling somewhere around a few percent. If we included some long bursts as part of the sample, as argued reasonable in Sec. V.2, the event rate would adjust further. Without a clear criterion for separating merger-driven and formed-in-isolation GRBs, it is very difficult to formally estimate precursor rarity and thus this remains an open problem in our view. The problem also goes the other way: is it strictly necessary to have a merger to form a short GRB [see 414]? This issue has also been discussed by Wang et al. 420 in light of the fact that even deducing $T_{90}$ is prone to systematics.
@@
The next standout problem is deducing the rarity of the subset of precursors that are launched premerger. It is likely that not all precursors are in fact launched pre-merger; if we took the GW/GRB 170817A waiting time of $\sim 1.7$ s as canonical, about $\sim 25\%$ of precursors would be associated with the inspiralling stars themselves (see Figure 14 below). The delay time is however rather sensitive to the post-merger environment. High-resolution, numerical studies show that black holes typically launch jets within $\lesssim 10^{2}$ ms while metastable neutron stars (if able at all; see Sec. V.7) tend to take longer [see 339, and references therein]. Black hole environments are typically cleaner as baryon pollutants from dynamical or post-merger ejecta, that could drag the jet and reduce its Lorentz factor, are swept under the event horizon, leading to shorter breakout timescales [see 299, for a discussion]. These issues are discussed in more detail in Sec. VI.1. Even so, combining the above, we may estimate that of order $\gtrsim 1\%$ binary mergers involving at least one neutron star display a premerger precursor.
@@ V.4 Spectral inferences
Spectral information about a precursor may be useful in reducing the uncertainties detailed above, as well as landing on a likely candidate for the ignition mechanism. As described in detail by Tsang et al. 402, Tsang 401, one may anticipate less thermality in cases where magnetic fields are dynamically dominant, at least with respect to premerger models. Essentially, the matter boils down to how important backreactions are. If the local magnetic fields in the region where the flare originates are strong, energy may be more freely transported along open field lines (via Alfvén waves; see also Sec. VI.5) without scattering that would otherwise tend to thermalise the spectrum. If instead the field is relatively weak there, cross-field transport is hardly prohibited and scattering can take place en masse. A typical decider in this matter will be how strong the magnetic energy density is compared to the precursor luminosity, with [402]
@@
In the above we have introduced the precursor luminosity, $L_{\rm prec}$. Deducing this quantity from observations is non-trivial, and few events in the literature have had luminosities reported. The difficulty stems from the fact that one needs to translate (i) photon counts into fluxes, and then (ii) fluxes into luminoisites. For the former this depends on the energy band of the instrument(s) and often requires delicate extrapolation between instruments [see 263, for a discussion on common pitfalls]. The latter requires knowledge of source redshift, often not recorded at the times of publishing because it requires hunting through astronomical databases for coincident events (there may not be any). One must also be careful in deducing a kind of averaged luminosity as opposed to a peak one, since the gamma-ray flux will, of course, be time-dependent. Thus far, no clear understanding of the precursor luminosity distribution is available in the literature [though see 111, for a recent study]. It should also be understood that $L_{\rm prec}$ refers to an isotropic luminosity, and thus could actually be lower if there is a non-negligible amount of beaming (see Sec. VI.3.1).
@@
Many bright precursor flares tend to exhibit a non-thermal spectrum [see table 2 in 443, for instance]. This would be expected if Alfvén waves propagating along open field lines are the primary means of the associated energy transport as detailed above, though this requires high field strengths and brings its own set of issues (cf. Sec. II.3). Regardless, brightness combined with non-thermality has led directly to the suggestion that some mergers contain magnetars [e.g. 398, 378, 225, 118, 428]. Precursor rarity may therefore be directly related to the prevalence of magnetars in mergers. Explaining how a strong field can persist into merger, even in $\lesssim 1\%$ of the population, remains an open issue under this interpretation. One possibility is that the field is not preserved at all but rather generated just before merger from convective instabilities triggered by mode-driven differential rotation; see Sec. VI.5.
@@
As some specific examples, Wang et al. 420 noted that the (bright) precursors in GRBs 111117A and 160804180 strongly favoured non-thermal fits. These both had short waiting times relative to the main event, and could thus favour a magnetospheric interaction or high-frequency mode resonance (see Sec. VI for a review of theoretical elements). Zhong et al. 443 noted a strong favouring for a non-thermal, cutoff power-law in the (dim) precursor from GRB 130310A, which occurred $\sim 4$ s prior to the main event. The precursors of GRBs 081216 and 141102A, both with delay times of $\sim 1$ s, can be convincingly fit with a blackbody [417]. They may therefore be associated with weak fields, or some kind of shock-breakout model. The absence of thermal components in the extremely bright precursor from GRB 211211A [$L_{\rm prec}\approx 7\times 10^{49}\text{ erg s}^{-1}$; 428], which occurred $\sim 1$ s prior to the main event could theoretically be expected from either a Poynting-dominated outflow from a remnant or possibly premerger magnetar interaction. Systematising such an analysis would constitute a useful addition to the literature, though spectral fittings can be notoriously complicated and this extends yet another observational problem therefore. Some recent efforts towards systematic, spectral classification were carried out by Deng et al. 111.
@@ V.5 Waiting times
It is evident that there is a significant spread of precursor phenomena, with a couple orders of magnitude separating the shortest and longest waiting times and also durations. It is probable therefore that there are multiple ignition mechanisms, involving either pre- or post-merger phenomena; the physics pertaining to delays are discussed in Secs. VI.1 and VI.3.3. We remark on the nature of some entries in Tab. 3, as some have error bars while others do not. In Wang et al. 420, from which many of our tabulated precursor results are quoted from, the authors did not given a confidence interval attached to the $\pm$ values they provided. As per standard convention, these error bars may correspond to $84.13\%$ confidence upper and lower limits containing the $68.27\%$ confidence interval [154]. In some other studies, such as Minaev and Pozanenko 270, Zhong et al. 443, waiting times are given as a kind of mean value from $T_{90}$ data (see Sec. V.1), which we quote here using the $\sim$ symbol. Owing to these differences in notation and convention, on top of other systematics described thus far, caution should be applied when weighting precursor significance across different studies.
Table 3: List of precursors (likely) associated with merger events, organised in descending order of waiting timing relative to main event, with some remarks on special features (if any). Key: XAM = X-ray afterglow with internal plateau, possibly suggestive of a magnetar remnant as identified by Rowlinson et al. 334 and others [e.g. 377, 379, 380]. Most data were compiled by Wang and Liu 417, though others not listed there come from Minaev and Pozanenko 270 (071030), Zhong et al. 443 (100717, 130310A, 170726794), Dichiara et al. 118 (230307A), and Deng et al. 111 (180703B). The long-form Fermi GBM catalogue name is used for some GRBs. Some precursors may be postmerger.
@@
071030 $0.9\pm 0.2$ $\sim 2.5$ $2.7\pm 0.5$ Data appear lost [see 443]? Debated $T_{90}$
@@ V.6.1 GRB 211211A
GRB 211211A was accompanied by a kilonova [323]. This by itself is exciting: the GRB was long $(T_{90}\gtrsim 30$s), and yet the association with a kilonova identifies the origin as a merger event (see Sec. V.2). Aside from this, the event showed a precursor flare $\sim 1$s prior to the main event [428]. In Suvorov et al. 380, the precursor timing was matched with a $g$-mode resonance with sufficiently large overlap that the crust may break, within the resonant failure model (see Sec. VI.3). Since magnetic fields skew the mode eigenfrequency, the magnetic field strength that was obtained $(\gtrsim 10^{14}$ G) was required to be consistent with that set by the extraordinarily high isotropic luminosity $(\sim 7\times 10^{49}\text{erg s}^{-1}$) and the absence of thermal emissions in the precursor (see Sec. V.4). Aside from this, the event was modulated by quasi-periodic oscillations (QPOs) at $\sim 22$ Hz [428] [see also 86].
The excitation of seismic modes in the crust is a natural prediction of a failure: a deep quake relieves energy from the crust which then causes the surroundings to vibrate through aftershocks. These aftershocks are most conducive to torsional modes, as initially pointed out by Duncan 129, which happen to have frequencies in the $\sim 20$ Hz range. The combination of having mass estimates for the binary members from the kilonova [323], estimates for the internal properties and magnetic field strength from a $g$-mode resonance, and estimates for the shear modulus in the crust from the torsional mode frequency proves intriguing. Although a thorough parameter space scan was not performed by Suvorov et al. 380, it was shown that conventional assumptions on the EOS (Fig. 2) and shear modulus from recent literature can accommodate all of the observed properties. More precisely, the frequency of the precursor QPO (interpreted as a torsional mode) and relative timing of the precursor itself (interpreted as a $g$-mode resonance) can be well-fit with a star that has kilonova-inferred mass $M=1.25M_{\odot}$ [323] with the APR EOS (Sec. II.1.2) and shear modulus computed from nuclear physics (Sec. III.1). Other interpretations are valid, though. For example, the QPO could be associated with a magnetoelastic mode [e.g. 149, 150] and the precursor could be attributed to an interface-mode [e.g. 402, 401, see also Sec. VI.4.1].
@@ V.6.2 GRB 180703B
The 180703B precursor is unusual because of the event duration. As found by Deng et al. 111, the precursor emissions began $\sim 1.8$ s prior to the main burst, lasting all the way until a lull at $\sim 0.3$ s, followed by the GRB of a similar duration. The delay could therefore be longer depending on what one calls the waiting time (again see Sec. V.1). The authors however denoted this precursor as a “gold sample” burst of very high significance with a thermally-dominated spectrum. The main GRB was also well fit with strongly thermal components and overall high-degree of spectral similarity. These facts together may point towards a common origin, i.e. a postmerger model with a choking jet or similar (see Sec. VI.6).
@@ V.6.4 GRB 191221A
The precursor from 191221A is currently the record holder for earliest precursor relative to main event, with a staggering $\sim 20$ s waiting time (though cf. Sec. V.2.1). No constraints on the spectrum were placed [420]. This event was coincident with LVK’s O3b run though no GWs were detected [9], pointing to a non-negligible source redshift. The distinct lack of any X-ray afterglow is also potentially problematic with respect to a stable neutron-star remnant scenario, which would probably be favoured in order to ascribe the long delay to a jet formation timescale [441, 73]. Premerger models in this direction, assuming a $<10$ s combined jet production and breakout timescale (see Sec. VI.1), are practically limited to low eigenfrequency resonances such as $g$-modes or ocean modes (Sec. VI.4) unless strong magnetic fields, thermal gradients, or rotation skew the eigenfrequency towards lower values (see Sec. IV.3). A systematic analysis of this event would be worthwhile to carry out because of the above considerations.

### v2

@@ V.2.1 Excluded events
The first two of these are described by Troja et al. 403, and are remarkable in that the waiting times exceed $\sim 100$ s. The authors remark several times however that they “are unable to confidently determine whether they are real features” as they were not seen by instruments other than Swift and are of relatively low significance. Although the possibility of such a long waiting time could offer exciting insight into neutron-star structure (cf. Fig. 9), we do not comment on these events further here. The final BATSE event is described by Koshut et al. 221, also with a very long waiting time of $\sim 75$ s. The significance is, however, similarly unclear (as is perhaps visually evident from the fourth-to-last panel seen in their figure 3). Note that Wang et al. 425 found a precursor from 080702A but much closer to the main event; it could be, therefore, that this is the second example of a known double precursor event after 090510 (see Sec. V.6.5). In the case of GRB 0911117, Troja et al. 403 find a significance of only $\sim 1.8\sigma$ and Zhong et al. 448 remark that the data can no longer be downloaded from the Swift archives (although this event may have also been recorded by Suzaku). We exclude this otherwise unremarkable event also (waiting time $\sim 2.7$ s).
@@ V.3 Rarity
Uncertainties notwithstanding with respect to the long-short classification issues detailed above, classifications for merger-driven precursors also depend on a variety of systematics. For instance, some authors require that a genuine precursor must precede the main burst by more than the burst $T_{90}$ [e.g. 403, 276]. This would mean there is a clear separation of timescales (as in Fig. 13), and not a possible blur period where it is unclear whether the precursor was related to burst ramp-up (cf. GRB 150922A). Some other studies allow for arbitrarily short waiting times relative to the main burst for early flashes to qualify for precursor status [see 448, 97, 425, for discussions].
@@
Another more obvious issue relates to sample size and instruments, with some events only being seen by Fermi, and/or Swift, and/or INTEGRAL [some events even by Agile, KONUS, Suzaku, Messenger, and HEND-Odyssey; 276, cf. Table 3]. As such one has to make a statistical decision about the significance required to define a precursor with respect to one or more instruments. Taking a high cut-off for some information criterion (e.g. Akaike or Bayesian) will naturally reduce the event rate and hence predict that precursors are rarer. As noted by Suvorov and Kokkotas 383 and others, even these deceptively small systematics can change event rate predictions from $\sim 0.4\%$ [276] to $\sim 10\%$ [403], with the most recent studies falling somewhere around a few percent. If we included some long bursts as part of the sample, as argued reasonable in Sec. V.2, the event rate would adjust further. Without a clear criterion for separating merger-driven and formed-in-isolation GRBs, it is very difficult to formally estimate precursor rarity and thus this remains an open problem in our view. The problem also goes the other way: is it strictly necessary to have a merger to form a short GRB [see 419]? This issue has also been discussed by Wang et al. 425 in light of the fact that even deducing $T_{90}$ is prone to systematics.
@@
The next standout problem is deducing the rarity of the subset of precursors that are launched premerger. It is likely that not all precursors are in fact launched pre-merger; if we took the GW/GRB 170817A waiting time of $\sim 1.7$ s as canonical, about $\sim 25\%$ of precursors would be associated with the inspiralling stars themselves (see Figure 14 below). The delay time is however rather sensitive to the post-merger environment. High-resolution, numerical studies show that black holes typically launch jets within $\lesssim 10^{2}$ ms while metastable neutron stars (if able at all; see Sec. V.7) tend to take longer [see 345, and references therein]. Black hole environments are typically cleaner as baryon pollutants from dynamical or post-merger ejecta, that could drag the jet and reduce its Lorentz factor, are swept under the event horizon, leading to shorter breakout timescales [see 305, for a discussion]. These issues are discussed in more detail in Sec. VI.1. Even so, combining the above, we may estimate that of order $\gtrsim 1\%$ binary mergers involving at least one neutron star display a premerger precursor.
@@ V.4 Spectral inferences
Spectral information about a precursor may be useful in reducing the uncertainties detailed above, as well as landing on a likely candidate for the ignition mechanism. As described in detail by Tsang et al. 407, Tsang 406, one may anticipate less thermality in cases where magnetic fields are dynamically dominant, at least with respect to premerger models. Essentially, the matter boils down to how important backreactions are. If the local magnetic fields in the region where the flare originates are strong, energy may be more freely transported along open field lines (via Alfvén waves; see also Sec. VI.5) without scattering that would otherwise tend to thermalise the spectrum. If instead the field is relatively weak there, cross-field transport is hardly prohibited and scattering can take place en masse. A typical decider in this matter will be how strong the magnetic energy density is compared to the precursor luminosity, with [407]
@@
In the above we have introduced the precursor luminosity, $L_{\rm prec}$. Deducing this quantity from observations is non-trivial, and few events in the literature have had luminosities reported. The difficulty stems from the fact that one needs to translate (i) photon counts into fluxes, and then (ii) fluxes into luminoisites. For the former this depends on the energy band of the instrument(s) and often requires delicate extrapolation between instruments [see 269, for a discussion on common pitfalls]. The latter requires knowledge of source redshift, often not recorded at the times of publishing because it requires hunting through astronomical databases for coincident events (there may not be any). One must also be careful in deducing a kind of averaged luminosity as opposed to a peak one, since the gamma-ray flux will, of course, be time-dependent. Thus far, no clear understanding of the precursor luminosity distribution is available in the literature [though see 112, for a recent study]. It should also be understood that $L_{\rm prec}$ refers to an isotropic luminosity, and thus could actually be lower if there is a non-negligible amount of beaming (see Sec. VI.3.1).
@@
Many bright precursor flares tend to exhibit a non-thermal spectrum [see table 2 in 448, for instance]. This would be expected if Alfvén waves propagating along open field lines are the primary means of the associated energy transport as detailed above, though this requires high field strengths and brings its own set of issues (cf. Sec. II.3). Regardless, brightness combined with non-thermality has led directly to the suggestion that some mergers contain magnetars [e.g. 403, 383, 231, 119, 433]. Precursor rarity may therefore be directly related to the prevalence of magnetars in mergers. Explaining how a strong field can persist into merger, even in $\lesssim 1\%$ of the population, remains an open issue under this interpretation. One possibility is that the field is not preserved at all but rather generated just before merger from convective instabilities triggered by mode-driven differential rotation; see Sec. VI.5.
@@
As some specific examples, Wang et al. 425 noted that the (bright) precursors in GRBs 111117A and 160804180 strongly favoured non-thermal fits. These both had short waiting times relative to the main event, and could thus favour a magnetospheric interaction or high-frequency mode resonance (see Sec. VI for a review of theoretical elements). Zhong et al. 448 noted a strong favouring for a non-thermal, cutoff power-law in the (dim) precursor from GRB 130310A, which occurred $\sim 4$ s prior to the main event. The precursors of GRBs 081216 and 141102A, both with delay times of $\sim 1$ s, can be convincingly fit with a blackbody [422]. They may therefore be associated with weak fields, or some kind of shock-breakout model. The absence of thermal components in the extremely bright precursor from GRB 211211A [$L_{\rm prec}\approx 7\times 10^{49}\text{ erg s}^{-1}$; 433], which occurred $\sim 1$ s prior to the main event could theoretically be expected from either a Poynting-dominated outflow from a remnant or possibly premerger magnetar interaction. Systematising such an analysis would constitute a useful addition to the literature, though spectral fittings can be notoriously complicated and this extends yet another observational problem therefore. Some recent efforts towards systematic, spectral classification were carried out by Deng et al. 112.
@@ V.5 Waiting times
It is evident that there is a significant spread of precursor phenomena, with a couple orders of magnitude separating the shortest and longest waiting times and also durations. It is probable therefore that there are multiple ignition mechanisms, involving either pre- or post-merger phenomena; the physics pertaining to delays are discussed in Secs. VI.1 and VI.3.3. We remark on the nature of some entries in Tab. 3, as some have error bars while others do not. In Wang et al. 425, from which many of our tabulated precursor results are quoted from, the authors did not given a confidence interval attached to the $\pm$ values they provided. As per standard convention, these error bars may correspond to $84.13\%$ confidence upper and lower limits containing the $68.27\%$ confidence interval [155]. In some other studies, such as Minaev and Pozanenko 276, Zhong et al. 448, waiting times are given as a kind of mean value from $T_{90}$ data (see Sec. V.1), which we quote here using the $\sim$ symbol. Owing to these differences in notation and convention, on top of other systematics described thus far, caution should be applied when weighting precursor significance across different studies.
Table 3: Precursors (likely) associated with merger events, organised in descending order of waiting timing relative to main event, with remarks on special features (if any). Key: XAM = X-ray afterglow with internal plateau, suggestive of a magnetar remnant as identified by Rowlinson et al. 340 and others [e.g. 384]. Most data were compiled by Wang and Liu 422, though others not listed there come from Minaev and Pozanenko 276 (071030), Zhong et al. 448 (100717, 130310A, 170726794), Dichiara et al. 119 (230307A), and Deng et al. 112 (180703B). The long-form Fermi GBM catalogue name is used for some GRBs. Some precursors may be postmerger.
@@
071030 $0.9\pm 0.2$ $\sim 2.5$ $2.7\pm 0.5$ Data appear lost [see 448]? Debated $T_{90}$
@@ V.6.1 GRB 211211A
GRB 211211A was accompanied by a kilonova [328]. This by itself is exciting: the GRB was long $(T_{90}\gtrsim 30$s), and yet the association with a kilonova identifies the origin as a merger event (see Sec. V.2). Aside from this, the event showed a precursor flare $\sim 1$s prior to the main event [433]. In Suvorov et al. 385, the precursor timing was matched with a $g$-mode resonance with sufficiently large overlap that the crust may break, within the resonant failure model (see Sec. VI.3). Since magnetic fields skew the mode eigenfrequency, the magnetic field strength that was obtained $(\gtrsim 10^{14}$ G) was required to be consistent with that set by the extraordinarily high isotropic luminosity $(\sim 7\times 10^{49}\text{erg s}^{-1}$) and the absence of thermal emissions in the precursor (see Sec. V.4). Aside from this, the event was modulated by quasi-periodic oscillations (QPOs) at $\sim 22$ Hz [433] [see also 87].
Seismic excitations in the crust is a natural prediction of a failure: a deep quake relieves energy from the crust which then causes the surroundings to vibrate through aftershocks. These aftershocks are most conducive to torsional modes, as initially pointed out by Duncan 130, which happen to have frequencies in the $\sim 20$ Hz range. The combination of having mass estimates for the binary members from the kilonova [328], estimates for the internal properties and magnetic field strength from a $g$-mode resonance, and estimates for the shear modulus in the crust from the torsional mode frequency proves intriguing. Although a thorough parameter space scan was not performed by Suvorov et al. 385, it was shown that conventional assumptions on the EOS (Fig. 2) and shear modulus from recent literature can accommodate all of the observed properties. More precisely, the frequency of the precursor QPO (interpreted as a torsional mode) and relative timing of the precursor itself (interpreted as a $g$-mode resonance) can be well-fit with a star that has kilonova-inferred mass $M=1.25M_{\odot}$ [328] with the APR EOS (Sec. II.1.2) and shear modulus computed from nuclear physics (Sec. III.1). Other interpretations are valid, though. For example, the QPO could be associated with a magnetoelastic mode [e.g. 150, 151] and the precursor could be attributed to an interface-mode [e.g. 407, 406, see also Sec. VI.4.1].
@@ V.6.2 GRB 180703B
The 180703B precursor is unusual because of the event duration. As found by Deng et al. 112, the precursor emissions began $\sim 1.8$ s prior to the main burst, lasting all the way until a lull at $\sim 0.3$ s, followed by the GRB of a similar duration. The delay could therefore be longer depending on what one calls the waiting time (again see Sec. V.1). The authors however denoted this precursor as a “gold sample” burst of very high significance with a thermally-dominated spectrum. The main GRB was also well fit with strongly thermal components and overall high-degree of spectral similarity. These facts together may point towards a common origin, i.e. a postmerger model with a choking jet or similar (see Sec. VI.6).
@@ V.6.4 GRB 191221A
The precursor from 191221A is currently the record holder for earliest precursor relative to main event, with a staggering $\sim 20$ s waiting time (though cf. Sec. V.2.1). No constraints on the spectrum were placed [425]. This event was coincident with LVK’s O3b run though no GWs were detected [9], pointing to a non-negligible source redshift. The distinct lack of any X-ray afterglow is also problematic with respect to a stable neutron-star remnant scenario, which would probably be favoured in order to ascribe the long delay to a jet formation timescale [446, 74]. Premerger models in this direction, assuming a $<10$ s combined jet production and breakout timescale (see Sec. VI.1), are practically limited to low eigenfrequency resonances such as $g$-modes or ocean modes (Sec. VI.4) unless strong magnetic fields, thermal gradients, or rotation skew the eigenfrequency towards lower values (see Sec. IV.3). A systematic analysis of this event would be worthwhile to carry out because of the above considerations.